# A new indeterminacy-based quantum theory

### v4

@@ Acknowledgments
Appendix A: The Wigner’s friend experiences
@@
One might object that in some extended Wigner’s friend theorems, [^77] See Bong et al. (2020). it’s plausible to consider that the friend Alice inside her isolated lab sees a determinate outcome. In a sense, this theorem assumes that Wigner, without performing any operations on Alice and her lab and after her measurement, simply opens her lab door and asks her about what outcome she obtained. In the simple case where the friends share a spin $1/2$ entangled particles, she will answer that she obtained spin-up or spin-down with $50\%$ of probability each (i.e., if Wigner makes a projective measurement on the state of Alice after her measurement, without performing any other operation on the lab, he will obtain these outcomes). So, it seems that Alice sees a determinate outcome contrary to what EnDQT claims in (the highly idealized) situations where we manage to isolate the friend from interacting with SDCs. To put the objection more dramatically, the measurement problem can be regarded as the problem of accounting for the experiences of determinate outcomes of experimentalists upon measurements, despite QT predicting that in measurement-like interaction can yield indeterminate outcomes. The friend inside the isolated lab seems to experience a determinate outcome, but EnDQT seems to give no account of what this agent is experiencing. Hence, EnDQT doesn’t solve the measurement problem.
@@
First, note that in the case where we manage to isolate the lab’s contents from the SDCs, according to EnDQT, Wigner opening the lab triggers a physical process that leads to Alice obtaining determinate outcomes and reporting them to Wigner. It’s not necessarily the case that Alice sees a determinate outcome inside her lab before opening the door. Seeing a determinate outcome can arise due to the interactions with the SDCs when the door is opened.
@@
Second, as I have mentioned in the main text, there are different positions one may adopt regarding the friend’s experiences, and which one is the correct one depends on deep philosophical and empirical issues, which I don’t have space to settle here. The main point that I want to make now is that the above objection is not worrying, and there are different ways of answering it. On top of that, I will argue that the above objection could also be applied to other more accepted interpretations of QT in certain circumstances (such as the MWI), and in so far, it is a legitimate worry, it could also be a worry applicable to these interpretations. Given how accepted these interpretations are, it shouldn’t be a reason to reject EnDQT.
@@
Regarding the different positions, as I have mentioned in the main text, a possible one is that a) the agent lacks mental content underlying its perception of the outcome: this is the absent experience hypothesis. The claim would be that we shouldn’t worry that EnDQT (and other interpretations, as we will see below) can lead us to friend-like agents without experiences. We shouldn’t follow our intuitions in the extreme (and quite possibly unrealistic) environments of a completely isolated agent and think that that agent will be exactly like us. The problem with this possibility is that it’s hard to make sense of an agent without mental content.
@@
However, as I have also mentioned, EnDQT can even consider that the friend experiences something in the isolated lab via particular hypotheses, dissolving the above worry. We might consider that b) friend-like systems in isolated regions have indeterminate mental content, where this content depends on the indeterminate physical properties of their brain. When the lab is open, their indeterminate mental content becomes determinate. I will call this possibility the quantum experience hypothesis. This hypothesis has the benefit of not being foreign to the philosophy of mind. Externalism about mental content is roughly the thesis that mental content depends on the external environment of the subject that has that mental content. [^78] Putnam (1975). The friend $having$ determinate mental content is dependent on an environment that renders that content determinate.
@@
To deal with these cases, together with b), we can adopt a version of the extended mind hypothesis of Clark & Chalmers (1998), which I will call the quantum extended mind hypothesis. The idea is that the bearers of the friend’s mental content would be the outputs of the friend to the exterior (i,e, the sheet of paper). At first, their mental content is indeterminate; then it becomes determinate when the box is opened. Like the most sophisticated technology is perhaps an extension of our mind, for an incredible agent like the friend, its outputs that interact with the external environment are an extension of their mind. Note that there isn’t any action at a distance here according to EnDQT. So, Wigner measuring the output doesn’t influence the friend’s body. [^79] Note also that the extended quantum mind thesis differs from the traditional extended mind thesis by considering that even phenomenal content can have extended bearers. I don’t see any problem with considering that. More concretely, some might justify the extended mind thesis via individuating mental content through its functional roles (Clark & Chalmers, 1998). However, some may reject the claim that phenomenal content can be individuated by its functional roles (e.g., Chalmers, 1996). It’s unclear that my thesis requires a functionalist account of phenomenal content. I will leave the investigation of this topic for future work.

### v5

@@ Acknowledgments
Appendix A: Toy-model of an SDC
In this section, I will present a non-relativistic toy model of an SDC. The notation of this model may look complicated, but I will essentially be modeling a chain of simple decohering interactions. Thus, I will also provide a simple and pedagogical example of decoherence.
Since I am considering non-relativistic quantum theory, it cannot be guaranteed that the laws will be Lorentz and, more generally, general covariant. However, I will provide indications that these symmetries will be respected in a relativistic model. These indications will come from showing that given CDC1)-CDC4), all successive events where systems have determinate values will be time-like separated from each other. So, the order of events, which is important for EnDQT to maintain (to explain how determinacy arises), won’t vary according to the reference frame, and the events that arise from SDCs arise locally (although the duration of interactions can be described according to different reference frames). Also, even if systems are entangled with each other, since I am not reifying the wavefunction if a system has a determinate value (i.e., “collapses”), it doesn’t affect any other system that was entangled with. I will get back to these points at the end of this section.
I will represent each system in the model by $s^{k}_{ij}$, and each system will occupy a fixed spatial point that I will specify further below in a way that will emulate a quantum field in a spatial point (modulo the difference in degrees of freedom). This will facilitate the transition to quantum field theory in future work. $i$ in $s^{k}_{ij}$ will represent a layer in the graph in which each system is. The higher the value of $i$, the later will be the events that the system $s^{k}_{ij}$ will give rise to (see Figure 1 with an example of a graph with three layers with systems with their indices). Thus, $i$ represents a temporal order. $k$ will represent groups of systems in each layer $i$. Each group of systems $k$ in a layer $i$ will serve as an $environment$ that will decohere a single system $s^{l}_{ij}$ in the layer $i+1$, where in $i+1$ that system will have an index $j=k$. So, the upper script $k$ will help us designate which groups of systems in a layer $i$ decohere which systems in a layer $i+1$. When I omit specific indices, I will be implicitly referring to all systems that have the indices that weren’t omitted. So, when I write $S_{0}\rightarrow S_{1}\rightarrow S_{2}$, I will be referring to all the interactions between all systems with $S_{i=0}$ (in the layer $i=0$) and $S_{i=1}$ (in the layer $i=1$), as well as the interactions between all systems with $S_{i=1}$ and systems with $S_{i=2}$, where these interactions obey the conditions CDC1)-CDC4).
Importantly, I will assume that the temporal order of interactions in the graphs runs from the top to the bottom: systems $S_{0}$ first interact with $S_{1}$, and then (in agreement with the CDCs), systems $S_{1}$ interact with $S_{2}$. As I have said, this can also be written as $S_{0}\rightarrow S_{1}\rightarrow S_{2}$, where the temporal order of interactions runs from left to right. The interaction between systems $S_{0}$ and $S_{1}$ gives rise to the set of spatiotemporal events that I will designate as $E_{S_{0}\rightarrow S_{1}}$, which corresponds to systems $S_{0}$ and $S_{1}$ having determinate values. The interaction between systems $S_{1}$ and $S_{2}$ gives rise to the set of events $E_{S_{1}\rightarrow S_{2}}$. Note that given CDC1), events $E_{S_{0}\rightarrow S_{1}}$ are time-like separated from the event $E_{S_{1}\rightarrow S_{2}}$, although some events in the set $E_{S_{0}\rightarrow S_{1}}$ may be space-like separated from each other. This gives rise to the following “causal” order of events that I will write like this $E_{S_{0}\rightarrow S_{1}}\rightarrow E_{S_{1}\rightarrow S_{2}}$, where the temporal order of events runs from the left to the right. So, each edge in these graphs corresponds to one event, which involves both relata of interactions having a determinate value.
Figure 7: Representation of an SDC with the the structure $S_{0}\rightarrow S_{1}\rightarrow S_{2}$ and with three labels. Temporal order runs from the top to the bottom. Each edge corresponds to one event, which involves both relata of interactions having a determinate value. In this graph, we have 12 interactions depicted, which corresponds to 12 events.
In agreement with postulate CDC1*-ii*), let’s consider that in this SDC, systems $S_{i}$ start interacting with systems $S_{i-1}$ while $S_{i-1}$ are $already$ interacting with systems $S_{i-2}$ (where $S_{i-2}$ have the DC-$S_{i-1}$) so that $S_{i-1}$ has the DC-$S_{i}$, and $S_{i-1}$ can end up transmitting the DC to $S_{i}$ concerning some other system that $S_{i}$ might end up interacting with. However, when $S_{i}$ and $S_{i-1}$ begin interacting, let’s assume that we can neglect the evolution of the quantum states of $S_{i-1}$ while $S_{i-2}$ and $S_{i-1}$ interact, such that we can consider that systems $S_{i-1}$ and $S_{i}$ start interacting only when the interaction between $S_{i-2}$ and $S_{i-1}$ ends. In a more complicated model, this assumption will be translated in terms of different timings in which the coupling constants $g$ in the Hamiltonians specified above are non-zero.123123123As I have explained in Section 2, there is a way of making this assumption more precise via the commutativity criterion.
For simplicity, I assume that each interaction between a pair of systems, $S_{i}\rightarrow S_{i+1}$, constitutes a step in the model. The first step involves the interaction between systems $S_{0}$ and $S_{1}$, the second step will involve systems $S_{1}$ and $S_{2}$, and so on. At each step, I implement a state update, which corresponds to one of the possible determinate values of the relata of interactions that indeterministically arise with a certain probability. The probability for these determinate values is predicted via the reduced density operator for each interaction (more on this below).
Let’s turn to the Hamiltonian for the decoherence models, which will represent the interactions between the systems $S_{i}$ and $S_{i+1}$. I will start with the total Hamiltonian for this chain. I omit the indices that represent the systems and the groups of systems within each layer. I will have $N+1$ number layers, where $S_{0}\rightarrow S_{1}\rightarrow\ldots\rightarrow S_{N}$, where $N$ is the value of the index of the systems in the last layer, and where I start counting the layers from $i=0$. I will adopt a certain reference frame to describe these interactions. For simplicity, I will assume that each pair of systems $S_{i}$ and $S_{i+1}$ interact in a given frame during a fixed duration $\Delta$ that corresponds to the duration of each step. To express this interaction succinctly via the Hamiltonian of the SDC, let’s define the function $\chi(t_{i}+i\Delta,\Delta)$ given by the following difference between Heaviside functions:124124124 $\theta(x)=\begin{cases}0&\text{if }x<0\\ 1&\text{if }x\geq 0\end{cases}$
$\chi(t,i\Delta,\Delta)=\theta(t-i\Delta)-\theta(t-(i\Delta+\Delta)).$ (16)
Now, I can express the sequential interaction between systems that obey the laws given by CDC1)-CDC4), where their interactions start at $t=0$,
$\hat{H}=\sum^{N}_{i=0}\hat{H}_{S_{i}\rightarrow S_{i+1}}\chi(t,i\Delta,\Delta).$ (17)
$\hat{H}_{S_{i}}\rightarrow S_{i+1}$ designates the Hamiltonians of interaction describing the interaction between systems $S_{i}$ and $S_{i+1}$. As we will see, due to certain approximations, I will neglect the self-Hamiltonians of all systems. $S_{0}$ and its subsystems will be the initiators, which means that they will have the DC by default. Thus, they are able to transmit the DC to any other systems without having to interact with other systems that have the DC, and, therefore, they are able to start an SDC.
Consider the following interactions between systems that belong to two layers, given by $S_{i}\rightarrow S_{i+1}$. As it was mentioned above, I am assuming that if we have a system $S_{i+1}$ with $a$ number of elementary subsystems, we will have a system $S_{i}$ with $N_{i}=a$ number of groups of systems, which will constitute the environment that will decohere each elementary subsystem of $S_{i+1}$. Thus, in a toy model with the chain $S_{0}\rightarrow S_{1}\rightarrow\ldots\rightarrow S_{N}$, the higher the index $i$ of $S_{i}$, the lower the number of elementary subsystems each $S_{i}$ needs to have for this model to be adequate in describing processes of decoherence.
I will examine interactions modeled by the well-known spin-spin decoherence model [109, 29] because it is the most straightforward non-relativistic model. So, I will not be concerned with solving the model in general since it has been solved elsewhere. Rather, I will solve it for the simplest case to show how we can model SDCs via this decoherence model.
The initiator $S_{0}$ is constituted by sets of (initiator) two-level/spin-$1/2$ systems. Each set of systems interacts with one two-level/spin-$\frac{1}{2}$ system that is a subsystem of a larger system $S_{1}$. Furthermore, the spin-spin model will represent the successive interactions between the non-initiator systems, $S_{1}\rightarrow S_{2}\rightarrow\ldots\rightarrow S_{N}$. I will consider that in these interactions, the timescales in which systems are affected by their interactions with other systems are much shorter than their intrinsic evolution, and thus I will ignore their self-Hamiltonians. For reasons that will become clearer, I will divide the Hamiltonian in eq. (18) into two Hamiltonians,
$\begin{split}\hat{H}&=\sum_{i=0,\text{i even}}^{i=N}\hat{H}_{S_{i}\rightarrow S_{i+1}}\chi(t,i\Delta,\Delta)\\ &\quad+\sum_{i=1,\text{i odd}}^{i=N}\hat{H}_{S_{i}\rightarrow S_{i+1}}\chi(t,i\Delta,\Delta).\end{split}$ (18)
Let’s consider $N^{k}_{i}$ as a variable whose values concern the numbers assigned to the last member of the group $k$ in a layer $i$. Let’s consider $N^{k^{\prime}}_{i}$ as a variable whose values concern the numbers assigned to the first members of the group $k$ in a layer $i$. The Hamiltonian of the kind presented below will be applied to all the interactions between subsystems of $S_{i}$ and $S_{i+1}$, i.e., $S_{i}\rightarrow S_{i+1}$, where $i$ is odd. So, they will correspond to a specific $i$, $k$, and $l$ where, as a reminder, $k$ will be the group number in the layer $i$ whose value corresponds to an elementary system in the layer $i+1$. $l$ is the group number of systems in the layer $i+1$:
$\hat{H}_{s^{k}_{i}\rightarrow s^{l}_{(i+1)k}}$ $=\frac{1}{2}\hat{\sigma}_{xs^{l}_{(i+1)k}}\otimes\sum_{j=N^{k^{\prime}}_{i}}^{j=N^{k}_{i}}g_{s_{ij}^{k}\to s_{(i+1)k}^{l}}\hat{\sigma}_{xs^{k}_{ij}}$ (19)
$=\frac{1}{2}\hat{\sigma}_{xs^{l}_{(i+1)k}}\otimes\hat{E}_{s^{k}_{i}}.$
This Hamiltonian obeys the commutativity criterion of the following kind,
$[\hat{H},\sigma_{x}]=0,$ (20)
and thus $\sigma_{x}$ will be the pointer observable selected by the interactions represented via this Hamiltonian (Section 2).
On the other hand, the Hamiltonian below will apply to specific $i$, $k$, and $l$ where $S_{i}\rightarrow S_{i+1}$ and $i$ is even:
$\hat{H}_{s^{k}_{i}\rightarrow s^{l}_{(i+1)k}}$ $=\frac{1}{2}\hat{\sigma}_{zs^{l}_{(i+1)k}}\otimes\sum_{j=N^{k^{\prime}}_{i}}^{j=N^{k}_{i}}g_{s_{ij}^{k}\to s_{(i+1)k}^{l}}\hat{\sigma}_{zs^{k}_{ij}}$ (21)
$=\frac{1}{2}\hat{\sigma}_{zs^{l}_{(i+1)k}}\otimes\hat{E}_{s^{k}_{i}}.$
In this case, this Hamiltonian obeys the commutativity criterion of the following kind,
$[\hat{H},\sigma_{z}]=0,$ (22)
and thus $\sigma_{z}$ will be the pointer observable selected by the interactions represented via this Hamiltonian.
Let $k^{tot}_{i}$ be a variable that designates the total number of groups within each layer $i$. I am thus assuming that the number $k^{tot}_{i}$ of groups of systems of each $S_{i}$ will be equal to the total number of systems that constitute $S_{i+1}$, where $S_{i}\rightarrow S_{i+1}$.
I will assume the following initial states of all the systems that will give rise to an SDC,
$|\Psi(t=0)\rangle$ $=|\psi\rangle_{S_{0}}\bigotimes^{i=N}_{i=1,\text{i odd}}|\psi\rangle_{S_{i}}\bigotimes^{i=N}_{i=2,\text{i even}}|\psi\rangle_{S_{i}},$ (23)
Then, we have that the initial states of the initiators are
$|\psi\rangle_{S_{0}}=\bigotimes_{j=N^{1^{\prime}}_{1}}^{N^{1}_{1}}|+\rangle_{s^{1}_{1j}}\bigotimes_{j=N^{k^{\prime}}_{1}}^{N^{k}_{1}}\ldots\bigotimes_{j=N^{k^{\prime\text{tot}}_{1}}}^{N^{k^{\text{tot}}_{1}}_{1}}|+\rangle_{s^{k^{\text{tot}}_{1}}_{1j}}.$ (24)
For the non-initiator, we have for $i$ odd,
$|\psi\rangle_{S_{i}}=\bigotimes_{j=N^{1^{\prime}}_{i}}^{N^{1}_{i}}|+\rangle_{s^{1}_{ij}}\bigotimes_{j=N^{k^{\prime}}_{i}}^{N^{k}_{i}}\ldots\bigotimes_{j=N^{k^{\prime\text{tot}}_{i}}}^{N^{k^{\text{tot}}_{i}}_{i}}|+\rangle_{s^{k^{\text{tot}}_{i}}_{ij}}.$ (25)
For the non-initiators, we have for $i$ even with $i>0$,
$|\psi\rangle_{S_{i}}=\bigotimes_{j=N^{1^{\prime}}_{i}}^{N^{1}_{i}}|0\rangle_{s^{1}_{ij}}\bigotimes_{j=N^{k^{\prime}}_{i}}^{N^{k}_{i}}\ldots\bigotimes_{j=N^{k^{\prime\text{tot}}_{i}}}^{N^{k^{\text{tot}}_{i}}_{i}}|0\rangle_{s^{k^{\text{tot}}_{i}}_{ij}}.$ (26)
These initial states are set up in such a way that, given the above Hamiltonians of interaction, there will be decoherence. These are the initial states of our toy universe, and I will consider that it is a brute fact why they are the way they are. So, I am postulating a kind of “past hypothesis” [1] that assumes that these are special initial states that are uncorrelated with each other.125125125More on this in Section 2.
Let’s turn to a more detailed analysis of the spin-spin interactions. I will focus on the case of the interactions $S_{i}\rightarrow S_{i+1}$ with $i$ even. The cases where $i$ is odd will be analogous and give rise to the same results, except that systems will have determinate values $+$ or --. First of all, it is important to notice that we can assume that the “environments” composed of non-initiator systems along the SDCs have initial states randomly distributed due to their previous interactions with other systems of an SDC. This is important because it allows us to consider that decoherence will end up occurring since the initial states of the environment need to be randomly distributed. Furthermore, we can assume that we have randomly distributed coupling constants since the strength of those interactions will be random, given the random initial states of environmental systems of non-initiator systems.126126126To see why, note that when it comes to overlap terms, we are summing over the different phases of the off-diagonal terms of the reduced density matrix of the target quantum system, and in order for the phases to cancel each other, they need to be randomly distributed. See, e.g., [89] for more details. The initial states of the non-initiator target systems and the initiator systems will have, as a brute fact, randomly distributed amplitudes. Thus, we have for specific values of $k$, $l$ and $i$, the following initial state:
$|\psi(0)\rangle_{s^{k}_{i}\rightarrow s^{l}_{(i+1)k}}=a|0\rangle_{s^{l}_{(i+1)k}}+b|1\rangle_{s^{l}_{(i+1)k}}\prod_{j=N^{k^{\prime}}_{i}}^{j=N^{k}_{i}}(\alpha_{s^{k}_{ij}}|0\rangle_{s^{k}_{ij}}+\beta_{s^{k}_{ij}}|1\rangle_{s^{k}_{ij}}).$ (27)
Afterward, we have that under the evolution driven by $\hat{H}_{s^{k}_{i}\rightarrow s^{l}_{(i+1)k}}$ in eq. (21),
$|\psi(t)\rangle_{s^{k}_{i}\rightarrow s^{l}_{(i+1)k}}=$ $\;a|0\rangle_{s^{l}_{(i+1)k}}\prod_{j=N^{k^{\prime}}_{i}}^{j=N^{k}_{i}}\left(\alpha_{s^{k}_{ij}}\exp(ig_{s^{k}_{ij}\rightarrow s^{l}_{(i+1)k}}t)|0\rangle_{s^{k}_{ij}}+\right.$ (28)
$\quad\left.\beta_{s^{k}_{ij}}\exp(-ig_{s^{k}_{ij}\rightarrow s^{l}_{(i+1)k}}t)|1\rangle_{s^{k}_{ij}}\right)$
$+b|1\rangle_{s^{l}_{(i+1)k}}\prod_{j=N^{k^{\prime}}_{i}}^{j=N^{k}_{i}}\left(\alpha_{s^{k}_{ij}}\exp(-ig_{s^{k}_{ij}\rightarrow s^{l}_{(i+1)k}}t)|0\rangle_{s^{k}_{ij}}+\right.$
$\quad\left.\beta_{s^{k}_{ij}}\exp(ig_{s^{k}_{ij}\rightarrow s^{l}_{(i+1)k}}t)|1\rangle_{s^{k}_{ij}}\right).$
When we trace out the degrees of freedom of the environmental systems, we obtain the following equation,
$\text{Tr}_{s^{k}_{i}}|\psi(t)\rangle_{s^{k}_{i}\rightarrow s^{l}_{(i+1)k}}\langle\psi(t)|=$ $\;|a|^{2}|0\rangle\langle 0|+z(t)ab^{*}|0\rangle\langle 1|$ (29)
$+z^{*}(t)a^{*}b|1\rangle\langle 0|+|b|^{2}|1\rangle\langle 1|,$
where it can be shown that127127127[109].
$z(t)=\prod_{j=N^{k^{\prime}}_{i}}^{j=N^{k}_{i}}\left[\cos 2g_{s^{k}_{ij}\rightarrow s^{l}_{(i+1)k}}t+i\left(|\alpha_{s^{k}_{ij}}|^{2}-|\beta_{s^{k}_{ij}}|^{2}\right)\sin 2g_{s^{k}_{ij}\rightarrow s^{l}_{(i+1)k}}t\right].$ (30)
We can make a simple analysis of the evolution of this expression for $|\alpha_{s^{k}_{ij}}|=|\beta_{s^{k}_{ij}}|$,
$z(t)=\prod_{j=N^{k^{\prime}}_{i}}^{j=N^{k}_{i}}\cos 2g_{s^{k}_{ij}\rightarrow s^{l}_{(i+1)k}}t.$ (31)
Then, we can analyze how many systems it takes for decoherence to occur for randomly distributed coupling constants $g_{s^{k}_{ij}\rightarrow s^{l}_{(i+1)k}}$, by analyzing when $z(t)$ goes quasi-irreversibly to zero over time.
(a) $G=6$
(b) $G=12$
(c) $G=17$
Figure 8: Plots for the decoherence factor $z(t)$ in eq. (31) with randomly distributed $g_{s^{k}_{ij}\rightarrow s^{l}_{(i+1)k}}$, taking values between 0 and 1 for different numbers $G$ of environmental systems. It is necessary for them to be randomly distributed in order for decoherence to occur. As we can see, the higher the value of $G$, the lower the fluctuations of the decoherence factor $z(t)$. Thus, the time it takes decoherence to occur can be estimated, for example, by directly looking at plots like these ones and seeing at what time $z(t)$ goes to zero on average without any further future oscillations.
More generally, [109, 29] have shown that for a broad kind of distribution of couplings $g$, and a sufficiently large amount of environmental systems, the function $z(t)$ exhibits over time an approximately Gaussian decay,
$z(t)\approx e^{-\Gamma^{2}t^{2}}.$ (32)
The exact value of the decay constant $\Gamma^{2}$ depends on the initial state of the environment and the distribution of the coupling constants $g_{i}$.
Let’s focus on the simple SDC model involving decoherence models that assume the simplification $|\alpha_{s^{k}_{ij}}|=|\beta_{s^{k}_{ij}}|$. This will be enough for our purposes of illustrating how an SDC can be modeled. Let’s consider that in this model, we have randomly distributed coupling constants $g$. Let’s designate the number of systems that constitute group $k$ in layer $i$ as $|N^{k}_{i}|$. To simplify, let’s assume that the groups of systems that constitute each $S_{i}$ will have the same number of subsystems, and that number will be equal to a constant $G$. Thus, I will designate by $|N^{k}_{i}|=G$ for all layers $i$ and groups $k$ within each layer. Thus, given decoherence, in order for this model to work, we will need that $G>>1$. I will make a more detailed analysis of this feature in Figure 8.
Since the results in Figure 8 hold across all the spin-spin models considered in this SDC model, we can see that systems will decohere each other, obeying CDC1)-CDC4), thus giving rise to determinate values and transmitting the DC. To see under which conditions this is obtained and how it is obtained in more detail, let’s see how many systems we need to have in the rest of the layers in order for systems in the last layer $N$ to have determinate values, where $N+1$ is the total number of layers. This layer will be the layer of systems $S_{N}$ with their respective subsystems. Given our simplifications, the following formula allows us to calculate the number of systems in layer $i$ that are needed in order for decoherence to occur in layer $N$, where $G$ needs to be an adequate number (such as a number higher than 17):
Number of systems in layer i $=\textit{Number of subsystems of $S_{N}$}$ (33)
×$\quad\times G^{\textit{layer number i}}.$
Given this simple model, the value of the index $N$ of the system in the last layer, $G$, and the number of systems in each layer, we can calculate the number of events involving systems having determinate values that arise in this SDC,
Number of events in an SDC $=\sum_{i=0}^{i=N-1}\text{Number of systems}$ (34)
in the layer $\quad\text{in the layer }N\times G^{i}.$
Note that systems have determinate values at interactions, so I will assume that they have a determinate value at the spatial location of the target system. Thus, we associate to each target system $s$ a spatial point $(x,y,z)$, and the environmental systems $E$ of $s$ and $s$ will have determinate value in the spatiotemporal point $(t,x,y,z)$. Note that I am simplifying here, and it is more appropriate that systems have values in spacetime regions. If we have 5 systems in layer $N$ and 30 systems per group, which is a reasonable number to give rise to decoherence (Figure 8), and the total of $4$ layers (and this $N=3$), the SDC will give rise to $139655$ events involving systems having determinate values.
Now, we can see how the dynamics of this SDC is going to develop by implementing the following algorithm via simple decoherence models like the one above (Figure 8). I will explain this algorithm schematically. Before implementing it, we should use eq. (34) to assign an appropriate number of systems to each layer, given the number of subsystems in $S_{N}$ that are assumed. Within each layer $i$, we should number each system with natural numbers $j$ with $j>1$, and group them in groups of size $G$. The last layer, $N$, just needs one group. As I have said, we make each group correspond to an individual system in the next layer by making the group number in the layer $i$ correspond to an individual system in the layer $i+1$ (or a subsystem of $S_{i+1}$). These numbers, $i$, $k$, and $l$ will individuate systems $s^{l}_{ik}$. At the end of each step, we update the states in agreement with the determinate values that will arise randomly with a given Born probability. Then, we have
-Even steps:
Step 0: We analyze a decoherence model with the initial states for the initiators given by eq. (24), which will be the subsystems of $S_{0}$ and will be the environmental systems. The target systems will be the subsystems of $S_{1}$ and will be given by eq. (26). The Hamiltonian for this interaction will be given by (21).
The rest of the even steps: We analyze decoherence models where the states |+ket|$\rangle$| + ⟩ and |ket|$\rangle$| - ⟩ of the systems that arise from the odd steps will serve as initial states of the environment and the states in eq. (25) will serve as the initial states of the target system. The Hamiltonian for this interaction will be given by eq. (21).
-Odd steps:
We analyze decoherence models where the states $|0\rangle$ and $|1\rangle$ of the systems that arise from the prior even step will serve as initial states of the environment and the states in eq. (26) will serve as the initial states of the target systems. The Hamiltonian that governs/describes these interactions will be given by eq. (19).
Since I have supposed that in a certain reference frame, the duration of the interaction $\Delta$ is the same in all the interactions, we get that the time in this frame for the systems in layer $N$ to have a determinate value is $\Delta\times N$, where we start counting time in our model at $t=0$. For simplicity, we can assume that $\Delta$ will be equal to the average decoherence time as measured in a rest frame.128128128And in a frame far from any strong gravitational field. Assuming these frames, allow us to consider that the decoherence timescale can only get longer under certain relativistic transformations. In the model with four layers, when $G=30$, after an analysis involving diverse random coupling constants, the average decoherence time is approximately $0.6$ seconds. So, it will take approximately $1.8$ seconds for the systems $S_{3}$ to have determinate values. As we increase $G$, this time becomes shorter.
I will end this section by mentioning some important features that this model illustrates. First, notice that I am making some important simplifications. Via the coupling constant $\chi$, I am assuming that systems in each layer start interacting at the same time in a given frame. In a more realistic model where this assumption is not made, the coupling constants $g$ for each elementary system would depend on time, and therefore, we would just have one kind of coupling constants $g$, instead of $\chi$ and $g$. Although this seems to give rise to a very complex model, as we will see in future work, when the more fundamental quantum field theoric models are taken into account, in principle, we just need to consider one kind of coupling constant mediating the interaction between fields. So, in principle, we can describe the above dynamics of the SDC using any other reference frame. However, only when we have a quantum field theoretic model of SDCs this reference frame independence will become manifest.
Second, note also that the duration of decoherence may vary between reference frames. However, given quantum field theory in curved spacetime,129129129[94]., this variation won’t be a problem as long as the laws are generally covariant.
Third, given the CDC1)-CDC4), the events involving systems in the layer $10,...,N-$ 1 that lead to the events involving systems in the layer $N$ (which necessarily involve the systems in layer $N-1$) having a determinate value are time-like separated from each other. So, they will assume a fixed temporal order in all reference frames. To see this, note that CDC1)-CDC4) establishes that in order for systems $S_{N}$ in layer $N$ to have a determinate value due to systems $S_{N-1}$ in layer $N-1$ at $t$ (where systems have values in the same spacetime point), systems $S_{N-1}$ in layer $N-1$ need to have a determinate value due to systems $S_{N-2}$ in layer $N-2$ at $t^{\prime}<t$ in a given reference frame, and so on for the rest of events involving interactions between systems in the other layers. Thus, given CDC1)-CDC4), all successive events where systems have determinate values will be time-like separated from each other. So, the order of events, which is important for EnDQT to maintain (since the criteria for systems to have determinate values depend on this order), won’t vary according to the reference frame.
Fourth, it should be kept in mind that since there is no literal collapse of the wavefunction but only a state update, even if the initial states of each system in each layer were initially entangled with all the other systems in that layer, they don’t cause the other systems to collapse. So, there is no action at a distance or the adoption of some preferred reference frame.
Fifth, the above model involves the simple spin-spin decoherence models, but in principle, any other decoherence model that involves local interactions can be used to model SDCs.
Appendix B: The Wigner’s friend experiences
@@
As I have mentioned in Section 3, one might object that in some extended Wigner’s friend theorems,130130130See [14]. it is plausible to consider that the friend Alice inside her isolated lab sees a determinate outcome. In a sense, this theorem assumes that Wigner, without performing any operations on Alice and her lab and after her measurement, simply opens her lab door and asks her about what outcome she obtained. In the simple case where the friends share a spin-$1/2$ entangled particles, she will answer that she obtained spin-up or spin-down with $50\%$ of probability each (i.e., if Wigner makes a projective measurement on the state of Alice after her measurement, without performing any other operation on the lab, he will obtain these outcomes). So, it seems that Alice sees a determinate outcome contrary to what EnDQT claims in (the highly idealized) situations where we manage to isolate the friend from interacting with SDCs. To put the objection more dramatically, the measurement problem can be regarded as the problem of accounting for the experiences of determinate outcomes of experimentalists upon measurements, despite QT predicting that measurement-like interactions can yield indeterminate outcomes. The friend inside the isolated lab seems to experience a determinate outcome, but EnDQT seems to give no account of what this agent is experiencing. Hence, EnDQT doesn’t solve the measurement problem.
@@
First, note that in the case where we manage to isolate the lab’s contents from the SDCs, according to EnDQT, Wigner opening the lab triggers a physical process that leads to Alice obtaining determinate outcomes and reporting them to Wigner. It is not necessarily the case that Alice sees a determinate outcome inside her lab before opening the door. Seeing a determinate outcome can arise due to the interactions with the SDCs when the door is opened.
@@
Second, there are different positions one may adopt regarding the friend’s experiences, and which one is the correct one depends on deep philosophical and empirical issues, which I don’t have space to settle here. The main point that I want to make now is that the above objection is not worrying, and there are different ways of answering it. On top of that, I will argue that the above objection could also be applied to other more accepted interpretations of QT in certain circumstances (such as the MWI), and in so far, it is a legitimate worry, it could also be a worry applicable to these interpretations. Given how accepted these interpretations are, it shouldn’t be a reason to reject EnDQT.
@@
Regarding the different positions, as I have mentioned in the main text, a possible one is that a) the agent lacks mental content underlying its perception of the outcome: this is the absent experience hypothesis. The claim would be that we shouldn’t worry that EnDQT (and other interpretations, as we will see below) can lead us to friend-like agents without experiences. We shouldn’t follow our intuitions in the extreme (and quite possibly unrealistic) environments of a completely isolated agent and think that that agent will be exactly like us. The problem with this possibility is that it is hard to make sense of an agent without mental content.
@@
However, as I have also mentioned, EnDQT can even consider that the friend experiences something in the isolated lab via particular hypotheses, dissolving the above worry. We might consider that b) friend-like systems in isolated regions have indeterminate mental content, where this content depends on the indeterminate physical properties of their brain. When the lab is open, their indeterminate mental content becomes determinate. I will call this possibility the quantum experience hypothesis. This hypothesis has the benefit of not being foreign to the philosophy of mind. Externalism about mental content is roughly the thesis that mental content depends on the external environment of the subject that has that mental content.131131131[84]. The friend $having$ determinate mental content is dependent on an environment that renders that content determinate.
@@
To deal with these cases, together with b), we can adopt a version of the extended mind hypothesis of [26], which I will call the quantum extended mind hypothesis. The idea is that the bearers of the friend’s mental content would be the outputs of the friend to the exterior (i,e, the sheet of paper). At first, their mental content is indeterminate; then it becomes determinate when the box is opened. Like the most sophisticated technology is perhaps an extension of our mind, for an agent like the friend, its outputs that interact with the external environment are an extension of their mind. Note that there isn’t any action at a distance here according to EnDQT. So, Wigner measuring the output doesn’t influence the friend’s body.132132132Note also that the extended quantum mind thesis differs from the traditional extended mind thesis by considering that even phenomenal content can have extended bearers. I don’t see any problem with considering that. More concretely, some might justify the extended mind thesis via individuating mental content through its functional roles [26]. However, some may reject the claim that phenomenal content can be individuated by its functional roles (e.g., [25]). It is unclear that my thesis requires a functionalist account of phenomenal content. I will leave the investigation of this topic for future work.